OpenStudy (anonymous):

s(t)=-6t^2+12t+24 find displacement over interval of [0,3]

OpenStudy (anonymous):

also find total distance travelled

OpenStudy (amistre64):

displacement: (3,s(3)) -(0,s(0)) --------- 3 ,s(3)-s(0) ; pythag \[\sqrt{3^2 +(s(3)-s(0))^2 }\] \[\sqrt{9 +(s(3)-s(0))^2 }\]

OpenStudy (amistre64):

distance traveled: \[\int_{0}^{3}\sqrt{1+(-6t^2+12t+24)'^2}dx\] \[\int_{0}^{3}\sqrt{1+(-12t+12)^2}dx\] \[\int_{0}^{3}\sqrt{1+(144t^2-288t +144)}dx\] \[\int_{0}^{3}\sqrt{1+144t^2-288t +144}\ dx\]

OpenStudy (amistre64):

im sure that simplifies inthe end :)

OpenStudy (anonymous):

yes..thanks

OpenStudy (anonymous):

i think you have to be careful here

OpenStudy (anonymous):

although i think the answer is easy enough to get. you can see from this equation that you have a parabola that faces down. it has a maximum at x = 1

OpenStudy (anonymous):

at \[x=0\] you are at 24 and at \[x=1\] you are at 30 so you have traveled a total distance of 6 then you turn around and go back

OpenStudy (anonymous):

at \[x=3\] you are at 6 if my arithmetic is correct. so from 30 to 6 is 24 and your total distance traveled is 6+ 24 = 30

OpenStudy (anonymous):

i am assuming here that \[s(t)\] is position. if it is speed, then i am totally wrong.

OpenStudy (anonymous):

but i think if it is just speed, then since your speed is positive over this entire interval, you can just take \[\int_0^3(-6t^2+12t+24)dt \] to get your answer